3. A car has a mass of 1600 kg. It has an engine which develops 35 kW when travelling at a speed of 70 km/h. Neglecting losses, determine the resistance to motion in N/kg.
1
Expert's answer
2020-12-14T07:29:57-0500

Answer

Power is given by

"P=Fv=(R-mg) v"

So resistance

"R=\frac{P}{v}+mg=\frac{35000}{70\times\frac{5}{18}}\\+1600\times9.8"

R=16580 N

So in newton per Kg

R="\frac{16580}{1600}=10.36N/kg"